SYMMETRY-ADAPTED TENSORIAL FORMALISM FOR THE SPECTROSCOPY OF THE SO_{2}F_{2} QUASI-SPHERICAL TOP: APPLICATION TO THE BENDING TRIAD

\underline{M. ROTGER}, V. BOUDON, M. LOETE, N. ZVEREVA-LOETE, Laboratoire de Physique de l’Université de Bourgogne, CNRS UMR 5027, 9, Avenue Alain Savary, BP 47870, F-21078 Dijon Cedex, France.

The techniques of symmetry-adapted tensorial formalism and of vibrational extrapolation developed since many years by the Dijon group have proved their efficiency for the spectroscopy of spherical-top molecules (CH_{4}, SF_{6}, …). We have extended these methods to the case of quasi-spherical tops such as SO_{2}F_{2}^{a}. This model has been used recently to perform the analysis of the ground state of this molecule^{b}. We present here a preliminary study concerning the analysis of the ν_{3}/ν_{7}/ν_{9} bending triad in the 550 cm^{−1} region. These results are compared to those obtained with the usual asymmetric-top approach^{c} . A set of programs for spectrum calculations and fits named C_{2v}TDS has been used and is freely available at the URL:

                               http://www.u-bourgogne.fr/LPUB/c2vTDS.html

aM. Rotger, V. Boudon and M. Loëte, J. Mol. Spectrosc., 216, 297-307, (2002).

bM. Rotger, V. Boudon, M. Loëte, L. Margulès, J. Demaison, H. Mäder, G. Winnewisser and H.S.P. Müller, J. Mol. Spectrosc., 222, 172-179, (2003).

cH. Bürger, J. Demaison, F. Hegelund, L. Margulès, I. Merke, J. Mol. Struct., 612, 133-141, (2002).